The standard form equation for a parabola with focus (2, –3) and directrix x = 5 is?
Answer by lwsshak3(11628)

The standard form equation for a parabola with focus (2, –3) and directrix x = 5 is?
***
Given data shows this is an equation of a parabola with horizontal axis of symmetry of the standard form: (y-k)^2=4p(x-h), (h,k) being the (x,y) coordinates of the vertex. Parabola opens leftwards.
..
For given equation:
x-coordinate of vertex: midway between focus and directrix on axis of symmetry, y=-3
Vertex:(7/2,-3)
p=distance from vertex to focus or directrix on axis of symmetry=3/2
4p=6
Equation:
(y+3)^2=-6(x-7/2)